Operations Report – March 15th

Operations Report 15-MAR-2020

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Saturday generator

Off at 6 am. At 7 pm, SOC was 99%. Generator was not run overnight.

Sunday Generator

6 am SOC was 76

9 am SOC was 67

1 pm SOC was 100

7 pm was 100

Generator will not be run overnight.

Diesel Reading – 60%

Station Propane Reading – 69 %

Ethanol Free Gasoline – 5.5 gallons

Water (loft tank): 40 gallons

Water Meter: n/a units

Water (static tank) – 300 gallons

Static to Loft Pump used – not known

Water in GreenHab – basically empty gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: no

Hours: 152.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Sojourner rover used: no

Hours: 104.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Spirit rover: no

Hours: 134.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: no

Hours: 70.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Curiosity rover used: no

Hours: 139.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: No

Reason for use: n/a

Oil Added? no

ATV Fuel Used: None

Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report.

HabCar used and why, where? To Hanksville to dump trash

CrewCar used and why, where? Not used, in GJ, we will need to go get it soon

General notes and comments: Nothing to report

Summary of internet: Nominal

Summary of suits and radios: All radios were removed from power,

turned off and batteries removed. Waiting to hear from Scott on what

I should do with the spacesuits.

Summary of Hab operations: Heat still off. Removed a bunch of things

that we left by crews. I also noticed that all of the plastic

mattress covered had been removed from the beds. I’m not sure when

that happened, but it means that those mattresses have not been

protected at all. The pillows are all really dirty and will be

removed.

Summary of GreenHab operations: Heater at 16 degrees. Door open

during the day.Watered all plants. Will be shutting down the GreenHab

in the next few days.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: I think the power loss overnight

was due to observatory operations. I hadn’t considered that in my

experiment to see if we need to keep the rental generator.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Beware the Ides of March!